# Large Model Empowered Streaming Semantic Communications for Speech Translation

### v1

@@ IV Numerical Results
The FLEURS [17] is adopted as the speech dataset to train and test LaSC-ST in the experiments. It covers speech utterances in over 102 languages. Without loss of generality, we consider speech in English, Chinese, and French. The edge server and local devices are deployed within the campus local area network of Imperial College London. The edge server consists of four NVIDIA H100 GPUs.
Algorithm 2 Testing algorithm of the LaSC-ST with dynamic speech segmentation mechanism.
1:Input: Batch of input speech $\boldsymbol{S}$, trained networks ${\mathfrak{T}}_{\mathrm{SC}}(\cdot)$, ${\mathfrak{T}}_{\mathrm{CE}}(\cdot)$, ${\mathfrak{T}}_{\mathrm{CD}}(\cdot)$, ${\mathfrak{T}}_{\mathrm{SP}}(\cdot)$, and larger speech model, fading channel $\boldsymbol{H}$ from testing channel set $\boldsymbol{\mathcal{H}}$, a wide range of testing SNR regime.
@@
10:              Upload $\boldsymbol{P}$ to the edge server and extract $\boldsymbol{F}$.
11:              Download $\boldsymbol{F}$ to the local device.
@@
15:              Upload $\widehat{\boldsymbol{F}}$ to the edge server and attain $\widehat{\boldsymbol{P}}$.
16:              Download $\widehat{\boldsymbol{F}}$ to the local device.
17:              ${\mathfrak{T}}_{\mathrm{SP}}(\widehat{\boldsymbol{F}})\rightarrow\widehat{\widetilde{\boldsymbol{S}}}_{t}$.
@@
The speech compressor consists of two CNNs. The channel encoder and decoder include three dense layers. Two CNNs and one dense layer are utilized in the speech predictor. The Meta Seamless Communication is invoked as the large speech model. The hyperparameters $p$ and $q$ are both set to 0.05. The thresholds $m$ and $n$ are 0.65 second and 0.85 second, respectively. The parameter settings of LaSC-ST are summarized in Table II.
TABLE I: Average latency of the translated speech segments predicted by different systems under Rayleigh channels.
Ground Truth DeepSC-S2T+VIST LaSC-ST (Fixed) LaSC-ST (Dynamic)
@@
TABLE II: Parameter settings of the proposed LaSC-ST.
Layer Name Parameters Activation
Speech Compressor 2$\times$1D CNN 128, 64 channels ReLU
Channel Encoder 3$\times$×Dense 3$\times$2048 units None
Channel Decoder 3$\times$×Dense 3$\times$2048 units ReLU
Speech Predictor 2$\times$1D CNN 2$\times$1280 channels ReLU
1$\times$×Dense 1 unit None
Large Speech Model Meta Seamless Communication
The BLASER 2.0 results are shown in Fig. 4, where the input language is English and the translated language is Chinese [^1] More simulations results of different languages and the translated speech samples can be found at https://github.com/Zhenzi-Weng/LaSC-ST.. A benchmark is provided by cascading a semantic communication system for speech-to-text translation, DeepSC-S2T [14], and a cutting-edge text-to-speech pipeline, VIST [18]. From the figure, the LaSC-ST outperforms the benchmark and attains BLASER 2.0 scores of over 3.0 in the high SNR regime, which verifies the effectiveness of the established edge-device collaborative semantic communication framework. Moreover, the LaSC-ST with the fixed speech segmentation approach provides superior quality of translated speech compared to the dynamic speech segmentation mechanism because the speech compressor and speech predictor are trained under the fixed-duration speech segments.
@@ V Conclusions
In this paper, we developed a large model-empowered semantic communication system to support streaming speech translation, named LaSC-ST. Particularly, the semantic extraction module is offloaded to the edge server to mitigate the computational demands on the local device. The large speech model is leveraged to break the language constraint of the input speech and generate unified semantic features. Moreover, a novel dynamic speech segmentation algorithm is proposed to reduce the end-to-end transmission latency by adaptively adjusting the speech segment duration.

### v2

@@ IV Numerical Results
The FLEURS [17] is adopted as the speech dataset to train and test LSSC-ST in the experiments. It covers speech utterances in 102 languages. Without loss of generality, we consider speech in English, Chinese, and French. The edge server and local devices are deployed within the campus local area network of Imperial College London. The edge server consists of six NVIDIA H100 GPUs.
The speech compressor consists of two CNNs. The channel encoder and decoder include three dense layers. Two CNNs and one dense layer are utilized in the speech predictor. The Meta Seamless Communication is invoked as the large speech model. Hyperparameters $p$ and $q$ are both set to 0.05. Thresholds $m$ and $n$ are 0.65 second and 0.85 second, respectively. The parameter settings of LSSC-ST are summarized in Table I.
Algorithm 2 Testing algorithm of the LSSC-ST with dynamic speech segmentation mechanism.
1:Input: Batch of input speech $\boldsymbol{S}$, trained networks ${\mathfrak{T}}_{\mathrm{SC}}(\cdot)$, ${\mathfrak{T}}_{\mathrm{CE}}(\cdot)$, ${\mathfrak{T}}_{\mathrm{CD}}(\cdot)$, ${\mathfrak{T}}_{\mathrm{SP}}(\cdot)$, and large speech model, fading channel $\boldsymbol{H}$ from testing channel set $\boldsymbol{\mathcal{H}}$, a wide range of testing SNR regime.
@@
10:              Upload $\boldsymbol{P}$ to the edge server.
11:              Extract $\boldsymbol{F}$ from $\boldsymbol{P}$.
@@
15:              Attain $\widehat{\boldsymbol{P}}$ from $\widehat{\boldsymbol{F}}$.
16:              Download $\widehat{\boldsymbol{P}}$ to the local device.
17:              ${\mathfrak{T}}_{\mathrm{SP}}(\widehat{\boldsymbol{P}})\rightarrow\widehat{\widetilde{\boldsymbol{S}}}_{t}$.
@@
The BLASER 2.0 results are shown in Fig. 4, where the input language is English and the translated language is Chinese [^1] More simulations results of different languages and the translated speech samples can be found at https://github.com/Zhenzi-Weng/LSSC-ST.. A benchmark is provided by cascading a semantic communication system for speech-to-text translation, DeepSC-S2T [18], and a cutting-edge text-to-speech pipeline, VIST [19]. From the figure, the LSSC-ST outperforms the benchmark and attains BLASER 2.0 scores of over 3.0 in the high SNR regime, which verifies the effectiveness of the established edge-device collaborative semantic communication framework. Moreover, the LSSC-ST with the fixed speech segmentation approach provides superior quality of translated speech compared to the dynamic speech segmentation mechanism because the speech compressor and speech predictor are trained under the fixed-duration speech segments.
TABLE II: Average latency of the translated speech segments predicted by different systems under Rayleigh channels.
Ground Truth DeepSC-S2T+VIST LSSC-ST (Fixed) LSSC-ST (Dynamic)
@@ V Conclusions
In this paper, we developed a large model-empowered semantic communication system to support streaming speech transmission, named LSSC-ST. Particularly, the semantic extraction and channel coding modules are offloaded to the edge server to mitigate the computational demands on the local device. The large speech model is leveraged to break the language constraint of the input speech, generating unified semantic features and supporting multilingual speech translation tasks. Moreover, a novel dynamic speech segmentation algorithm reduces the end-to-end transmission latency by adaptively adjusting the speech segment duration.